Two electromagnetic waves are traveling through a material. Wave 1 has a maximum electric field strength that is three times the maximum field strength of Wave 2. How do the average intensities of the waves compare?

a. The average intensity of Wave 1 is three times the intensity of Wave 2.

b. The average intensity of Wave 2 is three times the intensity of Wave 1.

c. The average intensity of Wave 1 is nine times the intensity of Wave 2.

d. The average intensity of Wave 2 is nine times the intensity of Wave 1.
Answer #1

, I = intensity . k = constant of proportionality E = Electric field amplitude

E1 = 3 x E2

(I1) = k^2 x (E1)^2 = k^2 x (3 E2)^2 =9 k^2 (E2)^2

(I2) = k^2 x (E2)^2

(I1)/ (i2) =9/1

( I1) = 9 x ( I2) . ( Answer C)